how do you remove brackets? e.g. 2(x+5) or a(a-4)


multiply out of brackets and simplify? e.g. (x+3)(+2) or (2c-3)(3c+4)

Answer by checkley71(8403)

2(X+5)=2X+10 ANSWER.
A(A-4)=A^2-4A ANSWER.
(X+3)(+2)=2X+6 ANSWER.
(2C-3)(3C+4)=6C^2-9C+8C-12=6C^2-C-12 ANSWER.